# Hyperradiance, photon blockade and concurrence in a pair of qubits inside a driven cavity

### v1

@@ Abstract
We theoretically study the radiance properties of a pair of qubits inside a single-mode cavity driven by a two-photon drive. Our results show that, when the two qubits are strongly coupled to the cavity field, the collective radiation emitted from the qubits exhibits hyperradiance which can be detected as a signature of two qubit entanglement in the weak-driving regime. We quantify the entanglement in terms of concurrence. Additionally, we study the radiance behaviour in the presence of an intracavity Kerr-nonlinear medium that leads to two-photon blockade. Our results suggest that this system with nonlinearity may act as a quadrature-squezzed and hyperradiant two-photon source.
@@ I Introduction
Cooperative effects of qubits [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11] interacting with a single-mode quantized field lead to the well-known phenomena of Dicke superradiance [12, 13, 14] and subradiance [12, 7, 5, 15, 16, 17]. Superradiance has been observed in collective emission from two-levels atoms [18, 19, 20, 21], nuclei [22, 23] and pair of ions [7]. In recent times, another cooperative effect called hyperradiance (HR) [24, 25, 26, 27, 28] wherein the emission rate surpasss the superradiant rate of emission has come to the fore. Apart from atomic or nuclear or any massive qubits, the massless photons in the context of cavity quantum electrodynamics (CQED) can also exhibit cooperative effects owing to some nonlinear processes. One such photonic cooperative effect is the photon blockade (PB) [29, 30, 30, 31, 32, 33, 34, 35, 36, 37] which forbids appearance of $(n+1)th$ photon in an $n$-photon blockade. It originates either from intrinsic nonlinearity associated with the dressed-state ladder in the strong-coupling CQED [29, 30, 30, 32, 33, 38] or from the interaction of cavity photons with an intracavity nonlinear medium [31, 36, 37]. Since a cavity field can mediate cooperative effects between a pair of atomic or ionic qubits inside a driven cavity, the question naturally arises if there is any interplay between the radiance and qubit-qubit entanglement or concurrence of the qubit-pair. Furthermore, if the cavity contains a nonlinear medium that can induce multi-photon blockade, then is there any connection of the photon blockade with the radiance and qubit-qubit entanglement? Concurrence is the most reliable measure of two-qubit entanglement, and has been used in the context of quantum cryptography [39], quantum teleportation [40] and generation of quantum random number [41]. But, concurrence has not yet been employed to explicate radiance, to the best of our knowledge.
@@
We quantify the radiance in terms of radiance witness $R$ defined as in Ref. [28]. The negativity of $R$ indicates subradiant regime implying that the radiation in the presence of two qubits is suppressed than that in the the presence of one qubit. The regime with $R=1$ is termed as superradiant regime. The collective radiation from the qubits characterised by $R>1$ belongs to hyperradiant regime. We demonstrate that, in our model $R$ can reach a value as high as 50 which is substantially higher than the previously reported value [24, 25, 26, 27, 28]. In order to reveal the squeezing property of the radiation we calculate the Wigner distribution [43] of the field. The fact that the Wigner distribution is found to be eliptical in the hyperradiant regime is a clear signature of the quadrature squeezing of the field. This quadrature squeezing is again confirmed by the klyshko’s criterion [44] and the even-odd oscillations of the photon number distributions. We also study the two-photon and three-photon Hanbury Brown-Twiss(HBT) type correlation functions. We find that both the correlations are greater than unity implying super-Poissonian photon statistics.
@@
The rest of this paper is organized as follows. In Sec. II we present our model and methods. In Sec.III we present and discuss our results. Finally, we conclude in Sec. IV.
@@ B Radiance and Concurrence
$R=\frac{\langle a^{\dagger}a\rangle_{2}-\sum_{i=1,2}\langle a^{\dagger}a\rangle_{1,i}}{\sum_{i=1,2}\langle a^{\dagger}a\rangle_{1,i}}$ (5)
@@
This involves the intracavity average photon number subject to the existence of one or two qubits inside the cavity. $\langle a^{\dagger}a\rangle_{2}$ denotes the average photon number when both the qubits are present in the cavity. The average photon number when only $i$th single qubit is present in the cavity is denoted by $\langle a^{\dagger}a\rangle_{1,i}$. $R=0$ implies that the radiance is emitted from uncorrelated qubits, whereas any finite value of $R$ indicates correlation between the qubits. The negativity of $R$ implies that the radiation is suppressed compared to that due to two qubits, corresponding to the so-called subradiance regime, while $R>0$ reveals that the radiation is enhanced due to the presence of both qubits. The value of $R=1$ specifically denotes that the radiation exhibits scaling proportional to the square of the number of qubits N ($\propto N^{2}$). This characteristic is termed superradiance. When qubits are confined within a cavity, they experience a backaction from the cavity field. This modifies their collective radiative behavior, opening up an intriguing new possibility of hyperradiance regime characterised by $R>1$.
To analyze this qubit-qubit entanglement we make use of the concurrence [50] defined by:
C Criteria for $n$ photon blockade ($n$PB)
[…]
$P_{n}\geq\Pi_{n},P_{m>n}<\Pi_{m>n}$ (7)
where $\Pi_{n}$ is the Poisson distribution given by $\Pi_{n}=\langle n\rangle^{n}\exp(-\langle n\rangle)/n!$. $\langle n\rangle$ is the average number of photons given by $\langle n\rangle=\rm{Tr}[a^{\dagger}a\rho]$. According to Eq. (7), the $n$PB effect increases the likelihood of $n$ photons while suppressing the probabilities of the photon numbers greater than $n$. The other criterion is based on $g^{(n)}(0)$, an equal-time nth-order correlation function. The mean photon number $\langle n\rangle$ in the case of weak driving is naturally very low, or $\langle n\rangle<<1$. The $n$PB effect must meet the following requirements
$g^{(n)}(0)\geq 1,g^{(n+1)}(0)<1$ (8)
meaning that the field has $(n+1)$th order sub-Poissonian photon statistics and the $n$th order poissonian or super-Poissonian photon statistics. In particular, the 2PB effect satisfies the correlation functions $g^{(2)}(0)\geq 1$ and $g^{(3)}(0)<1$. The 1PB effect can be characterized by $g^{(2)}(0)<<1$.
D Nonclassicality of the field
In order to seek possible connection between $n$-photon blockade and the nonclassical nature of the cavity field, we calculate three quantities namely squeezing, Klyshko’s number and radiance.
where $X_{\theta}=(ae^{-i\theta}+a^{\dagger}e^{i\theta})$ is the linear combination of the hermitian quadrature operators $X=\frac{(a+a^{\dagger})}{\sqrt{2}}$ and $Y=\frac{i(a^{\dagger}-a)}{\sqrt{2}}$. Any negative value of $S_{\theta}$ indicates the squeezing nature of the field characterized by narrowing of the quadrature distribution compared to that of a coherent state. To verify the nonclassicality of the photon number distribution, we use the Klyshko’s criterion [44]
$K_{n}=\frac{(n+1)P_{n-1}P_{n+1}}{nP_{n}^{2}}$ (10)
E Two-qubit dressed-state picture
In order to explain radiance-concurrence-photon blockade connection, it may be instructive to work in dressed-state picture. In the absence of coherent drive and decay processes, we develop the dressed-state structure for the system in order to have some insight of the physical processes taking place in the system . Throughout our calculations, we have taken the coupling ($g_{1}=g_{2}=g$) between the atoms, unlike in Ref.[30, 25, 30] where atomic drive creates pathways that interfere destructively. We have defined the collective two-qubit basis states as $|gg\rangle$,|±$|\pm\rangle=\frac{|ge>\pm|eg\rangle}{\sqrt{2}}$ and $|ee\rangle$ to describe the dynamics in the dressed-state picture. In the single-photon space the eigenvalues and eigenvectors (dressed states) are obtained as $|\psi_{0}^{(1)}\rangle=-|+,0\rangle$ with $E_{0}^{(1)}=0$ and $|\psi^{(1)}_{\pm}\rangle=\frac{|gg,1\rangle\pm|+,0\rangle}{\sqrt{2}}$ with $E_{\pm}^{(1)}=\pm\sqrt{2}g$. In the two-photon space, the eigenvectors are $|\psi^{(2)}_{0}\rangle=-\frac{1}{\sqrt{3}}|gg,2\rangle+\frac{\sqrt{2}}{\sqrt{3}}|ee,0\rangle$ and $|\phi^{(2)}_{0}\rangle=-|+,1\rangle$ with corresponding eigenvalue $E_{0}^{(2)}=0$. The another eigenvector is $\psi_{\pm}^{(2)}=\frac{1}{\sqrt{3}}|gg,2\rangle\pm\frac{1}{\sqrt{2}}|+,1\rangle+\frac{1}{\sqrt{6}}|ee,0\rangle$ with corresponding eigenvalue $E_{\pm}^{(2)}=\sqrt{6}g$. Due to the symmetric coupling ($g_{1}=g_{2}=g$) the antisymmetric Dicke state $|-,n\rangle$ is decoupled in the $n$photon space ($n=1,2$). In order to know the allowed transitions between the dressed states in the presence of a two photon dive, we calculate the transition dipole matrix elements between the dressed states as shown in Fig. 2. In the one-photon space, Kerr nonlinearity ($\chi$) will not affect the dressed states. It will act on the states with photon numbers $n\geq 2$. So, the one photon dressed states and dressed energies will be same as mentioned above i.e: without $\chi$. In the two photon space the hamiltonian in the basis $|gg,2\rangle,|\pm,1\rangle$ and $|ee,0\rangle$ will be,
The dressed energies and dressed states will be obtained by diagonalizing this Hamiltonian. we have plotted the dressed energies of the system in presence of $\chi$ in Fig. 3. We have set the parameters as $\Delta=0,\chi=6\gamma,10\gamma$. The dressed eigenenergies and dressed eigenstates in this particular parameter regime with $\chi=6,g=10\gamma$ are $E_{0}^{(2)}=-21.35,E_{1}^{(2)}=0,E_{2}^{(2)}=3.80,E_{3}^{(2)}=29.55$ and
$|\psi_{1}^{(2)}\rangle$ $=$ $-|+,1\rangle$
III Results and discussions
We first present the results in the absense of Kerr medium in subsection III(A) and then discuss on the effects of Kerr nonlinearity ($\chi$)in subsection III(B). All the frequency quantities are scaled by the qubit decay constant $\gamma$.
In this section we explore radiance, concurrence between two qubits, squeezing and nonclassicality of the field.But we find no photon blockade effect when $\chi=0$. Moreover, we find a connection between the radiance which is the property of the field and the concurrence that is a measure of entanglement between the qubits.
The radiance witness $R$ is plotted as a function of scaled detuning $\Delta/\gamma$ and driving strength $\eta/\gamma$ in Fig. 4 (a). Near $\Delta=0,\pm\sqrt{6}g/2\gamma$ the system exhibits a hyperradiant behaviour due to the multiphoton excitations to the dressed eigenstates as illustrated in the dressed-state level diagram shown in Fig. 2. The strengths of the transition $\psi^{(0)}\rightarrow\psi_{0,\pm}^{(2)}$ are same and it is $\sqrt{\frac{2}{3}}\eta$. So, unlike [25, 30] where only $\psi^{(0)}\rightarrow\psi_{0}^{(2)}$ transition was allowed, we have three distinct hyperradiance peaks at three different detunings. In fact, we find regimes where $R\approx 50$ which is greater than the previously reported results so far [28, 26, 27, 25, 51], to the best of our knowledge. To gain some analytical insight about the hyperradiance, we also address the problem with complex hamiltonian approach as discussed in the Appendix A and perturbatively obtain an expression of $R$ given by

### v3

@@ Abstract
We theoretically study the radiance properties of a pair of qubits inside a single-mode cavity driven by a two-photon drive. Our results show that, when the two qubits are strongly coupled to the cavity field, the collective radiation emitted from the qubits exhibits hyperradiance which can be detected as a signature of two qubit entanglement in the weak-driving regime. We quantify the entanglement in terms of concurrence. Additionally, we study the radiance behaviour in the presence of an intracavity Kerr-nonlinear medium that leads to two-photon blockade. Our results suggest that this system with nonlinearity may act as a quadrature-squeezed and hyperradiant two-photon source.
@@ I Introduction
Cooperative effects of qubits interacting with a single-mode quantized field lead to the well-known phenomena of Dicke superradiance [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14] and subradiance [15, 15, 7, 5, 16, 17, 18]. Superradiance has been observed in collective emission from two-levels atoms [19, 20, 21], nuclei [22, 23] and pair of ions [24, 7]. In recent times, another cooperative effect called hyperradiance (HR) [25, 26, 27, 28, 29] wherein the emission rate surpasss the superradiant rate of emission has come to the fore. Apart from atomic or nuclear or any massive qubits, the massless photons in the context of cavity quantum electrodynamics (CQED) can also exhibit cooperative effects owing to some nonlinear processes. One such photonic cooperative effect is the photon blockade (PB) [30, 31, 31, 32, 33, 34, 35, 36, 37, 38] which forbids appearance of $(n+1)th$ photon in an $n$-photon blockade. It originates either from intrinsic nonlinearity associated with the dressed-state ladder in the strong-coupling CQED [30, 31, 31, 33, 34, 39] or from the interaction of cavity photons with an intracavity nonlinear medium [32, 37, 38, 36, 40, 41, 42]. Since a cavity field can mediate cooperative effects between a pair of atomic or ionic qubits inside a driven cavity, the question naturally arises if there is any interplay between the radiance and qubit-qubit entanglement or concurrence of the qubit-pair. Furthermore, if the cavity contains a nonlinear medium that can induce multi-photon blockade, then is there any connection of the photon blockade with the radiance and qubit-qubit entanglement? Concurrence is the most reliable measure of two-qubit entanglement, and has been used in the context of quantum cryptography [43], quantum teleportation [44] and generation of quantum random number [45]. But, concurrence has not yet been employed to explicate radiance, to the best of our knowledge.
@@
We quantify the radiance in terms of radiance witness $R$ defined as in Ref. [28]. The negativity of $R$ indicates subradiant regime implying that the radiation in the presence of two qubits is suppressed in comparison to that in the the presence of one qubit. The regime with $R=1$ is termed as superradiant regime. The collective radiation from the qubits characterised by $R>1$ belongs to hyperradiant regime. We demonstrate that, in our model $R$ can reach a value as high as 50 which is substantially higher than the previously reported value [25, 26, 27, 28, 29]. In order to reveal the squeezing property of the radiation we calculate the Wigner distribution [47] of the field. The fact that the Wigner distribution is found to be eliptical in the hyperradiant regime is a clear signature of the quadrature squeezing of the field. This quadrature squeezing is again confirmed by the klyshko’s criterion [48] and the even-odd oscillations of the photon number distributions. We also study the two-photon and three-photon Hanbury Brown-Twiss(HBT) type correlation functions. We find that both the correlations are greater than unity implying super-Poissonian photon statistics.
@@
The rest of this paper is organized as follows: In Sec. II we present our model and methods. In Sec.III we present and discuss our results. Finally, we conclude in Sec. IV.
@@ A Hamiltonian
Figure 1: Schematic diagram of a pair of identical qubits inside a cavity that may contain a Kerr nonlinear medium. Here, we have shown a pair of two-level atom as qubits. The cavity is driven by a two-photon drive with strength $\eta$ and frequency $\omega_{d}$. Atoms have the transition frequency $\omega_{a}$. The $i$-th atom is coupled to the cavity with strength $g_{i}$ and $\gamma_{i}$ is its decay rate. The decay rate of the cavity is $\kappa$.
@@ B Radiance and Concurrence
$R=\frac{\langle a^{\dagger}a\rangle_{2}-2\langle a^{\dagger}a\rangle_{1}}{2\langle a^{\dagger}a\rangle_{1}}$ (5)
@@
This involves the intracavity average photon number subject to the existence of one or two qubits inside the cavity. $\langle a^{\dagger}a\rangle_{2}$ denotes the average photon number when both the qubits are present in the cavity. The average photon number when only $i$th single qubit is present in the cavity is denoted by $\langle a^{\dagger}a\rangle_{1}$. $R=0$ implies that the radiation is emitted from uncorrelated qubits, whereas any finite value of $R$ indicates the existance of a correlation between the qubits. The negativity of $R$ implies that the radiation is suppressed compared to that due to two qubits, corresponding to the so-called subradiance regime, while $R>0$ reveals that the radiation is enhanced due to the presence of both qubits. The value of $R=1$ specifically implies that the radiation exhibits scaling proportional to the square of the number of qubits as a characteristic feature of superradiance. When qubits are confined within a cavity, they experience a backaction from the cavity field. This modifies their collective radiative behavior, opening up an intriguing new possibility of hyperradiance regime characterised by $R>1$. We numerically study the properties of radiance solving the master Eq. (3). Furthermore, to gain some analytical insight about hyperradiance, we address the problem by wave function approach with a complex Hamiltonian in the weak driving regime as discussed in Appendix A. In this approach, the mean photon numbers for single- and two-qubit-cavity system in steady-state are given by
$\langle{\hat{a}}^{\dagger}a\rangle_{1}$ $=$ $|C_{g,2}|^{2}+2|C_{e,1}|^{2}$
$\langle{\hat{a}}^{\dagger}a\rangle_{2}$ $=$ $|C_{gg,1}|^{2}+2|C_{gg,2}|^{2},$ (6)
respectively, where $C_{gg,n},C_{g,2},C_{e,1}$ are the probability amplitudes of the states $|gg,n\rangle,|g,2\rangle,|e,1\rangle$ as defined in Appendix A. We thus obtain a perturbative expression of $R$ given by